OpenStudy (anonymous):

How do you put this into reduced row echelon form? $$ \left[\begin{array}{rrr|r} 3 & -2 & -1 & -8 \\ 5 & 1 & 3 & 23 \\ 4 & 1 & -5 & -18 \end{array}\right] $$

OpenStudy (anonymous):

1) first of all divide with 3 the whole 1st row to get pivot point 1 (all other elements below it should be zero ) . 2) then subtract 4 times of row 1 from 3rd row 3) now subtract 5 times of row one from 2nd row do this and let me know that you get .

OpenStudy (anonymous):

@apple_pi

OpenStudy (anonymous):

ok thanks i'll work on it

OpenStudy (anonymous):

there are steps after this to carried out .. if you want to completely do this first do the above 3 steps and then will let you the further steps..

OpenStudy (anonymous):

So now I have $$ \left[\begin{array}{rrr|r} 1 & \frac{-2}{3} & \frac{-1}{3} & \frac{-8}{3} \\ 0 & \frac{13}{3} & \frac{14}{3} & \frac{109}{3} \\ 0 & \frac{11}{3} & \frac{-11}{3} & \frac{-46}{3} \end{array}\right] $$

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):

1) ok now multiply the second row with 13/3 to get 1 at the first point 2) then subtract 11/3 times of 2nd row from the third row .

OpenStudy (anonymous):

1) multiply? would that possibly be divide by 13/3?

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):

i should have written to multiply with 3/13 ... yes if you divide with 13/3 same .. end result will be 1

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Oh, yeah I got it from here. Thanks

OpenStudy (anonymous):

let me give you online interactive tool for free to convert matrix to reduced ech form. go here http://www.math.odu.edu/~bogacki/cgi-bin/lat.cgi select Transforming a matrix to reduced row echelon form (third one in list) then select 3 rows and three columns then enter the left hand part of matrix it will teach you complete steps !